Here's my question:
One day, the temperature at 8:00 A.M. was 41 degrees Farenheight, and by 1:00 P.M. the temperature was 61 degrees Farenheight. What was the hourly rate of temperature change??
I got 4 degrees farenheight/hour. Thanks!!

Answer by stanbon(75887)

One day, the temperature at 8:00 A.M. was 41 degrees Farenheight, and by 1:00 P.M. the temperature was 61 degrees Farenheight. What was the hourly rate of temperature change??
I got 4 degrees farenheight/hour.
-------------------------
You have two points: (8,41) and (13,61)
slope = (61-41)/(13-8) = 20/5 = 2 degrees/hr.
=====================
Cheers,
Stan H.